AB INITIO STRUCTURES FOR 90^{∘}-TWISTED s-trans-1,3-BUTADIENE AND FOR CYCLOOCTATE-TRENE; THE NAKED sp^{2}-sp^{2} BOND AND OTHER DISCLOSURES ABOUT EQUILBRIUM STRUCTURES OF BUTADIENE

DAVID FELLER, Department of Chemistry, Washington State University, Pullman, WA 99164-4630; \underline{NORMAN C. CRAIG} and ALBERT R. MATLIN, Department of Chemistry and Biochemistry, Oberlin College, Oberlin, OH 44074.

The bond length of a carbon-carbon sp^{2}-sp^{2} σ-bond without the perturbing effects of π-electron interactions has been estimated by high-level ab initio calculations [CCSD(T)/cc-pVNQ with CBS extrapolation] on two prototypical systems: a 90^{∘}-twisted form of butadiene and the tub conformer of cyclooctatetraene. The butadiene system yields a value of 1.4818 Å, considerably longer than previous estimates. The corresponding bond length in cyclooctatetraene is slightly shorter due to some π-electron delocalization. Ab initio results for related equilibrium structures of other rotamers/transition states of butadiene along the internal rotation coordinate (C-C bond) will be reported and compared with an experimental potential function.^{a}

aR. Engeln, D. Consalvo, J. Reuss, Chem. Phys. 160, 427 (1992).